Find that solution of the system such that yı(0) = 1, 42(0) = 1. y1 = y1 + y2 + sin() y = 2y1 + cos(t)
Answer #1

9 = 4, + Z2 & sind 9, 10) = 1 y, = 2y + Cost Y, (o) = 1 » (0-1), - Y = sint o -24, +dy = cost 0 7 DID-1)Y, -DI2 = cost -84,